OpenStudy (anonymous):

Let f be an odd function, that is f(-x) =-f(x). We know that integral from 0^5 f(x) = 3. Then using properties of definite integrals we conclude that...

OpenStudy (anonymous):

the second equation is like this\[\int\limits_{0}^{5}fx=3\]

OpenStudy (anonymous):

is it pretty simple? just the zero replaces with -5 and -3 instead of 3?

OpenStudy (psymon):

There's a bunch of assumptions you can make really. But if it is an odd function, then: \[\int\limits_{-a}^{a}f(x)dx = 0\] So.....I suppose you could conclude: \[\int\limits_{-5}^{0}f(x)dx = -3\]

OpenStudy (anonymous):

ok but thankfully this a actually a multiple choice question so all my options have 5 on the top and -5 on the bottom of the integral sign

OpenStudy (psymon):

Ah. Yeah, if a function is odd then it's as I said: \[\int\limits_{-a}^{a}f(x)dx = 0\] Might as well throw it out there, but if a function is even then: \[\int\limits_{-a}^{a}f(x)dx = 2\int\limits_{0}^{a}f(x)dx\]